# When Indemnity Insurance Fails: Parametric Coverage under Binding Budget and Risk Constraints

### v4

@@ Abstract
In high-risk environments, traditional indemnity insurance is often unaffordable or ineffective, despite its well-known optimality under expected utility. We compare excess-of-loss indemnity insurance with parametric insurance within a common mean-variance framework, allowing for fixed costs, heterogeneous premium loadings, and binding budget constraints. We show that, once these realistic frictions are introduced, parametric insurance can yield higher welfare for risk-averse individuals, even under the same utility objective and without relying on behavioral assumptions. The welfare advantage arises precisely when indemnity insurance becomes impractical, and disappears once both contracts are unconstrained. Our results help reconcile classical insurance theory with the growing use of parametric risk transfer in high-risk settings.
@@ 1 Introduction
Insurance markets are increasingly strained in regions exposed to low-frequency, high-severity risks such as floods, wildfires, and cyclones. In many such areas, indemnity insurance has become prohibitively expensive or altogether unavailable, as rising hazard intensity, capital requirements, and fixed costs push premiums and deductibles to levels that leave households effectively uninsured. A key practical limitation is that household indemnity insurance is typically written on a full-value (sum-insured) basis: the premium is charged to cover the entire exposure, and the main lever to restore affordability is to raise the deductible. In high-risk areas, this mechanism can push deductibles to economically unrealistic levels (e.g., tens or hundreds of thousands of dollars), rendering the contract theoretically available but impractical, and delivering negligible effective protection. By contrast, parametric insurance is inherently scalable: households can purchase a smaller, explicitly bounded layer of protection (e.g., a fixed payment per event) without having to insure the full loss. Crucially, parametric coverage scales by adding first-dollar protection, whereas scaling indemnity insurance requires removing first-dollar coverage through higher deductibles, quickly rendering partial indemnity ineffective.
@@
Recent developments in flood and wildfire insurance markets illustrate this (un)insurability trend: insurers have withdrawn from parts of the California residential market, flood insurance take-up remains persistently low in high-risk zones, and coverage gaps are expanding even in advanced economies (e.g., Kousky and Cooke, 2012; Jarzabkowski et al., 2023; UN Office for Disaster Risk Reduction, 2025). Yet when disasters occur, losses are rarely borne privately alone. Instead, they are often socialized ex post through government relief, reconstruction grants, and implicit guarantees. This growing disconnect between private insurance provision and public risk-bearing raises a fundamental economic question: when full indemnification is no longer feasible, what forms of risk transfer can still improve individual welfare?
@@
Classical insurance theory provides a clear benchmark. Under broad and well-understood conditions, an excess-of-loss indemnity contract maximizes the expected utility of a risk-averse agent when premiums are loaded proportionally to expected losses (Arrow, 1971; Raviv, 1979). This result has shaped both theoretical and practical views of optimal insurance design for decades. However, it relies on assumptions that are increasingly violated in high-risk environments. In particular, it abstracts from differences in risk borne and fixed costs borne by insurers across contract types, as well as binding budget constraints faced by households. When insurance is expensive, the deductible that is still affordable becomes so large that indemnity insurance offers little or no effective protection, even though it remains optimal in a purely formal sense. In such cases, the classical optimality result continues to hold mathematically, but loses much of its economic content.
@@
This paper revisits the optimality of indemnity insurance in such settings by comparing it with parametric insurance. Parametric contracts replace loss-based indemnification with a fixed payment triggered by an objective index, such as the occurrence of a disaster event. While this design introduces basis risk (arising from imperfect loss matching), it eliminates most of the loss adjustment costs, allows for rapid payouts, and materially reduces the risk borne by insurers. As a result, parametric insurance can be priced with lower loadings and smaller fixed costs, and can remain feasible in environments where indemnity insurance breaks down. Despite its growing use in sovereign and corporate risk transfer, parametric insurance for households remains underdeveloped. That said, policy and industry discussions increasingly contemplate parametric household covers for catastrophic perils.
@@
Nevertheless, indemnity insurance is still widely considered the best (or only) way to provide effective protection. We show that this view is incomplete. Using a tractable model of frequency and severity, we compare indemnity and parametric insurance under the same expected mean–variance utility criterion. Losses arrive according to a Poisson process and have censored exponential severity. This specification can capture the salient features of natural disaster risk (low-frequency, high-severity) while allowing for closed-form solutions. The indemnity contract is of excess-of-loss type with a deductible, while the parametric contract pays a fixed amount per event and shares the same trigger, so there is no basis risk on frequency. Importantly, we allow for distinct premium loadings and fixed costs across contract types, reflecting the higher capital requirements, claims management costs, and reconstruction delays associated with indemnity insurance. At the sovereign level, parametric instruments are widely used as part of fiscal layering strategies, where rapid, rules-based payouts provide immediate liquidity after disasters while larger losses are addressed through slower mechanisms. A similar logic applies at the household level: even limited parametric payments can play a critical role in enabling early recovery and reducing reliance on ad hoc public assistance.
@@
These findings contribute to several strands of the literature. First, they complement the classical optimal insurance results by identifying conditions under which their practical relevance breaks down, without abandoning expected-utility-based evaluation. Our analysis complements the classical insurance literature following Arrow (1963) and Raviv (1979), which establishes the optimality of deductible insurance under proportional pricing and frictionless markets. Rather than challenging these results, we show how their economic relevance may erode when heterogeneous fixed costs and loadings, as well as budget constraints, are introduced. In this sense, the paper bridges the gap between foundational insurance theory and the observed expansion of parametric risk transfer in high-risk environments. Second, they add to the emerging economic analysis of parametric insurance, which has focused largely on sovereign and corporate applications (Clarke, 2016; Huang and Shi, 2024). Third, they speak directly to current policy debates on disaster risk financing, insurance affordability, and the appropriate role of governments in supporting risk transfer in high-risk areas. Our analysis suggests that parametric insurance should not be viewed merely as an inferior substitute for indemnity insurance, but as a potentially welfare-improving instrument precisely where traditional insurance fails.
@@ 2.1 Losses and benefits
Huang and Shi (2024, in a mean-variance setting similar to ours) compare indemnity insurance with index insurance. They show that improving the index reduces basis risk and brings the parametric cover closer to indemnity in expected-utility terms. Their results, therefore, support an important implication for our analysis: moving beyond the constant benefit $k$ and reducing basis risk through index triggers or richer indemnity functions can make the welfare case for parametric insurance even more compelling.
@@ 2.4 Research question: indemnity vs parametric insurance
The key economic motivation is that indemnity insurance is typically priced and designed to cover (or at least credibly underwrite) the full underlying loss distribution. When full cover becomes unaffordable, the main actuarial lever to reduce the premium is to increase the deductible—often to levels that are difficult to interpret as meaningful household protection (a mechanism we will illustrate numerically). By contrast, parametric insurance is naturally scalable: the insured can purchase a modest layer of per-event protection $k$ at a correspondingly modest premium, and the benefit can be paid quickly upon verification of the trigger, without requiring the household to incur a huge out-of-pocket loss before receiving any payment.
@@ 5.5 Implications for insurance regulation
A more flexible regulatory approach would recognize parametric insurance as a distinct class of risk-transfer instruments rather than a second-best substitute for indemnity. This includes clarifying disclosure requirements around basis risk, ensuring contract transparency, and facilitating experimentation with hybrid designs that combine parametric triggers and partial indemnification (Huang and Shi, 2024).

### v5

@@ Abstract
In high-risk environments, traditional indemnity insurance is often unaffordable or ineffective, despite its well-known optimality under expected utility. We compare excess-of-loss indemnity insurance with parametric insurance within a common mean-variance framework, allowing for fixed costs, heterogeneous premium loadings, and binding budget constraints. Motivated by the disaster insurance and risk-sharing literature, we show that, once these realistic frictions are introduced, parametric insurance can yield higher welfare for risk-averse individuals, even under the same utility objective and without relying on behavioral assumptions. The welfare advantage arises precisely when indemnity insurance becomes impractical, and disappears once both contracts are unconstrained. Our results help reconcile classical insurance theory with the growing use of parametric risk transfer in high-risk settings.
@@ 1 Introduction
Insurance markets are increasingly strained in regions exposed to low-frequency, high-severity risks such as floods, wildfires, and cyclones. In many such areas, indemnity insurance has become prohibitively expensive or altogether unavailable, as rising hazard intensity, capital requirements, and fixed costs push premiums and deductibles to levels that leave households effectively uninsured, (e.g., Kousky, 2019; Netusil et al., 2021; Swiss Re Institute, 2025). A key practical limitation is that household indemnity insurance is typically written on a full-value (sum-insured) basis: the premium is charged to cover the entire exposure, and the main lever to restore affordability is to raise the deductible. In high-risk areas, this mechanism can push deductibles to economically (and practically) unrealistic levels (e.g., tens or hundreds of thousands of dollars), rendering the contract theoretically available but impractical, and delivering negligible effective protection. By contrast, parametric insurance is inherently scalable: households can purchase a smaller, explicitly bounded layer of protection (e.g., a fixed payment per event) without having to insure the full loss. Crucially, parametric coverage scales by adding first-dollar protection, whereas scaling indemnity insurance requires removing first-dollar coverage through higher deductibles, quickly rendering partial indemnity ineffective.
@@
Recent developments in flood and wildfire insurance markets illustrate this (un)insurability trend: insurers have withdrawn from parts of the California residential market, flood insurance take-up remains persistently low in high-risk zones, and coverage gaps are expanding even in advanced economies (e.g., Kousky and Cooke, 2012; Jarzabkowski et al., 2023; UN Office for Disaster Risk Reduction, 2025). Yet when disasters occur, losses are rarely borne privately alone. Instead, they are often socialized ex post through government relief, reconstruction grants, and implicit guarantees, a pattern that has been widely documented in the disaster insurance literature (Kousky and Kunreuther, 2014; Kousky, 2018). This growing disconnect between private insurance provision and public risk-bearing raises a fundamental economic question: when full indemnification is no longer feasible, what forms of risk transfer can still improve individual welfare?
@@
Classical insurance theory provides a clear benchmark. Under broad and well-understood conditions, an excess-of-loss indemnity contract maximizes the expected utility of a risk-averse agent when premiums are loaded proportionally to expected losses (Arrow, 1971; Raviv, 1979). This result has shaped both theoretical and practical views of optimal insurance design for decades. However, it relies on assumptions that are increasingly violated in high-risk environments, as seen in Kousky and Cooke (2012); Froot (2001). In particular, it abstracts from differences in risk borne and fixed costs borne by insurers across contract types, as well as binding budget constraints faced by households. When insurance is expensive, the deductible that is still affordable becomes so large that indemnity insurance offers little or no effective protection, even though it remains optimal in a purely formal sense. In such cases, the classical optimality result continues to hold mathematically, but loses much of its economic content.
@@
This paper revisits the optimality of indemnity insurance from an implementation and affordability perspective, by comparing it with parametric insurance. Parametric contracts replace loss-based indemnification with a fixed payment triggered by an objective index, such as the occurrence of a disaster event. While this design introduces basis risk (arising from imperfect loss matching), it eliminates most of the loss adjustment costs, allows for rapid payouts, and materially reduces the risk borne by insurers (Kousky, 2019; Figueiredo et al., 2018; Kunreuther et al., 2013; Lin and Kwon, 2020). As a result, parametric insurance can be priced with lower loadings and smaller fixed costs, and can remain feasible in environments where indemnity insurance breaks down. Despite its growing use in sovereign and corporate risk transfer, parametric insurance for households remains limited in scale and adoption, particularly relative to the extensive theoretical and policy literature on index and disaster insurance (Clarke, 2016; Jensen et al., 2016; Carter et al., 2017). That said, policy and industry discussions increasingly contemplate parametric household covers for catastrophic perils.
A growing literature has analyzed index and parametric insurance, particularly in the context of agricultural risk, sovereign disaster finance, and developing economies (e.g., Clarke, 2016; Clarke and Dercon, 2016; Huang and Shi, 2024a; Kousky and Cooke, 2012). The present paper emphasizes basis risk, demand frictions, and the role of parametric instruments in providing partial protection and liquidity when indemnity insurance is unavailable or inefficient. Our paper is complementary to this literature, but distinct in focus and method. Rather than studying index quality, take-up, or behavioral responses, we ask a specific and unresolved theoretical question: when standard indemnity insurance remains formally optimal under classical criteria but becomes economically irrelevant due to fixed costs and budget constraints, can parametric insurance dominate under the same utility objective?
@@
Nevertheless, indemnity insurance is still widely considered the benchmark for effective protection. We show that this benchmark can lose practical relevance once affordability and implementation frictions are taken into account. Using a tractable model of frequency and severity, we compare indemnity and parametric insurance under the same expected mean–variance utility criterion. Losses arrive according to a Poisson process and have censored exponential severity. This specification can capture the salient features of natural disaster risk (low-frequency, high-severity) while allowing for closed-form solutions. The indemnity contract is of excess-of-loss type with a deductible, while the parametric contract pays a fixed amount per event and shares the same trigger, so there is no basis risk on frequency. Importantly, we allow for distinct premium loadings and fixed costs across contract types, reflecting the higher capital requirements, claims management costs, and reconstruction delays associated with indemnity insurance. At the sovereign level, parametric instruments are widely used as part of fiscal layering strategies, where rapid, rules-based payouts provide immediate liquidity after disasters while larger losses are addressed through slower mechanisms. A similar logic applies at the household level: even limited parametric payments can play a critical role in enabling early recovery and reducing reliance on ad hoc public assistance (e.g., Jarzabkowski et al., 2024).
@@
These findings contribute to several strands of the literature.
Firstly, they complement the classical optimal insurance results by identifying conditions under which their practical relevance breaks down, without abandoning expected-utility-based evaluation. Our analysis complements the classical insurance literature following Arrow (1963) and Raviv (1979), which establishes the optimality of deductible insurance under proportional pricing and frictionless markets. Subsequent work shows that the structure of optimal indemnity insurance is preserved under more general environments, including non-linear premium principles and broader contract constraints (Huberman et al., 1983; Blazenko, 1985; Gollier and Schlesinger, 1996). Rather than challenging these results, we show how their economic relevance may erode when heterogeneous fixed costs and loadings, as well as budget constraints, are introduced. In this sense, the paper bridges the gap between foundational insurance theory and the observed expansion of parametric risk transfer in high-risk environments.
Secondly, they add to the economic analysis of parametric insurance. Existing work has examined the demand for index insurance and its welfare implications, particularly in agricultural and low-income settings (e.g., Clarke, 2016; Jensen et al., 2016; Carter et al., 2017; Chantarat et al., 2017), as well as the role of basis risk and index quality in contract performance and how improved trigger design can mitigate discrepancies between actual losses and index-based payouts (Cummins et al., 2004; Teh and Woolnough, 2019; Zhang et al., 2019; Clement et al., 2018). By contrast, our analysis abstracts from index optimization and treats the trigger structure as given, focusing instead on how affordability constraints and implementation frictions shape the relative performance of parametric and indemnity insurance. A related strand emphasizes the use of parametric insurance for extreme risks and catastrophe coverage, highlighting challenges of solvency, tail risk, and capital constraints that limit the feasibility of full indemnification in high-risk environments (Clement et al., 2018; Lopez and Thomas, 2023). The literature has largely focused on sovereign and corporate applications (Clarke, 2016; Huang and Shi, 2024b).
Thirdly, they speak directly to current policy debates on disaster risk financing, insurance affordability, and the appropriate role of governments in supporting risk transfer in high-risk areas. Our analysis suggests that parametric insurance should not be viewed merely as an inferior substitute for indemnity insurance, but as a potentially welfare-improving instrument precisely where traditional insurance fails. While recent work has compared parametric and indemnity insurance in environments with informational frictions, solvency constraints, or general equilibrium interactions (Feess et al., 2022; Wei, 2024; Lopez and Nkameni, 2025), we show that parametric insurance can outperform indemnity insurance even in the absence of informational imperfections or equilibrium feedbacks.
@@
To our knowledge, existing economic analyses of parametric insurance do not provide a closed-form welfare comparison with optimal excess-of-loss indemnity insurance under a common objective, nor do they isolate the role of fixed costs and budget constraints in rendering deductible insurance economically ineffective despite its classical optimality. By keeping preferences, pricing principles, and triggers aligned across contract types, our framework isolates a simple but underappreciated mechanism: parametric insurance remains scalable and welfare-improving precisely because it adds first-dollar protection, whereas partial indemnity insurance removes it.
@@ 2.1 Losses and benefits
Huang and Shi (2024b, in a mean-variance setting similar to ours) compare indemnity insurance with index insurance. They show that improving the index reduces basis risk and brings the parametric cover closer to indemnity in expected-utility terms. Their results, therefore, support an important implication for our analysis: moving beyond the constant benefit $k$ and reducing basis risk through index triggers or richer indemnity functions can make the welfare case for parametric insurance even more compelling.
@@ 2.4 Research question: indemnity vs parametric insurance
The key economic motivation is that indemnity insurance is typically priced and designed to cover (or at least credibly underwrite) the full underlying loss distribution. When full cover becomes unaffordable, the main actuarial lever to reduce the premium is to increase the deductible—often to levels that are difficult to interpret as meaningful household protection (a mechanism we will illustrate numerically). By contrast, parametric insurance is naturally scalable: the insured can purchase a modest layer of per-event protection $k$ at a correspondingly modest premium, and the benefit can be paid quickly upon verification of the trigger, without requiring the household to incur a huge out-of-pocket loss before receiving any payment. Unlike much of the existing literature on index insurance, which takes partial coverage as given, we explicitly compare how partial coverage is achieved under indemnity versus parametric designs, and show that the two scaling mechanisms have fundamentally different welfare implications.
@@ 5.5 Implications for insurance regulation
A more flexible regulatory approach would recognize parametric insurance as a distinct class of risk-transfer instruments rather than a second-best substitute for indemnity. This includes clarifying disclosure requirements around basis risk, ensuring contract transparency, and facilitating experimentation with hybrid designs that combine parametric triggers and partial indemnification (Huang and Shi, 2024b).